OpenStudy (anonymous):

Factor the polynomial r^4-k^4

OpenStudy (zehanz):

Hey @hogie! That one looks alot like the other! Hint: r^4 and k^4 are both squares, (of r² and k²), so you can use a²-b²=(a+b)(a-b) again. (and then another time...)

OpenStudy (anonymous):

=(r^2-k^2) (r^2+k^2) =(r-k)(r+k) [ (r+k)^2 -2rk)

OpenStudy (anonymous):